### v1

# The Mixtures and the Neural Critics: On the Pointwise Mutual Information Profiles of Fine Distributions

Abstract
Mutual information quantifies the dependence between two random variables and remains invariant under diffeomorphisms. In this paper, we explore the pointwise mutual information profile, an extension of mutual information that maintains this invariance. We analytically describe the profiles of multivariate normal distributions and introduce the family of fine distributions, for which the profile can be accurately approximated using Monte Carlo methods. We then show how fine distributions can be used to study the limitations of existing mutual information estimators, investigate the behavior of neural critics used in variational estimators, and understand the effect of experimental outliers on mutual information estimation. Finally, we show how fine distributions can be used to obtain model-based Bayesian estimates of mutual information, suitable for problems with available domain expertise in which uncertainty quantification is necessary.
1   Introduction
Mutual information (MI) is a non-parametric statistical measure used to determine the dependency between two random variables (r.v.) and numerous approaches have been proposed to estimate it (Kay, 1992; Kraskov et al., 2004; Cellucci et al., 2005; Belghazi et al., 2018). In recent work, Czyż et al. (2023) proposed a family of distributions for which ground-truth MI is analytically tractable to evaluate the performance of selected estimators on different distributions. They concluded that there is no universally acclaimed state-of-the-art estimator, with different methods performing best under different assumptions on the data-generating process, which had been often implicitly assumed. In parametric statistics, generative models are used to explicitly state the assumptions about the data-generating process. Although such model-based estimators have been used in settings with discrete r.v. (Hutter, 2001; Brillinger, 2004), the only model-based MI estimator for continuous r.v. is canonical correlation analysis (CCA), proposed by Kay (1992) and Brillinger (2004). As CCA assumes that the joint distribution is multivariate normal, we consider the development of more advanced model-based MI estimators to be an important open problem.
A second and related problem is the construction of new distributions with known ground-truth MI, which can be used to assess the performance of different estimators under varying distributional assumptions. Czyż et al. (2023) used multivariate normal and Student distributions (and their transformations) with analytically tractable MI to model the sparsity of interactions and effects of long tails. However, these distributions are far from universal: they are not suitable to model communication channels in which experimental failures result in measurement outliers or distributions used in biology (Grabowski et al., 2019) and physics (Carrara & Ernst, 2023), which involve both continuous and discrete r.v. Moreover, reparametrizations with diffeomorphisms do not change the pointwise mutual information profile (the distribution of pointwise mutual information), limiting the expressivity of constructed distributions.
In this work, we address both problems: we propose a family of fine distributions for which the ground-truth MI can be reliably estimated using Monte Carlo methods with arbitrarily high accuracy. The proposed distributions include normalizing flows (Kobyzev et al., 2021; Papamakarios et al., 2021) and their finite mixtures, allow for discrete and continuous random variables, and can be used to investigate the phenomena involving outliers resulting from failed experimental measurements. Moreover, we show how they can be used to create expressive generative models which provide model-based mutual information estimates, beyond the linear case of CCA.
Contributions
We introduce a novel family of fine distributions, for which mutual information and the pointwise mutual information profile can be accurately determined using Monte Carlo methods (Sec. 2). In Sec. 3 we demonstrate three different applications of fine distributions: first, we show how to construct low-dimensional distributions posing significant challenges to existing mutual information estimators (Sec. 3.1). Second, we use the introduced family of fine distributions to study the robustness of mutual information estimators to inliers and outliers (Sec. 3.2). Third, we show how to apply fine distributions to understand the biases and sample efficiency of variational estimator employing neural critics (Sec. 3.3). Finally, we demonstrate how the introduced family of  fine distributions can be used to provide model-based mutual information estimates (Sec. 4), allowing for Bayesian inference of mutual information.
Figure 1: First two panels: samples from a bivariate normal distribution and a transformed distribution. Both distributions have the same PMI profile (blue histogram in the fourth panel). Third panel: mixture distribution with distinct PMI profile, which cannot be obtained as a transformation of multivariate normal distribution due to a different PMI profile (pink histogram in the fourth panel). All three distributions have the same mutual information, marked with the black line in the fourth panel.
2   Theoretical framework
In this paper, we consider random variables valued in smooth manifolds without boundary (Lee, 2012, Ch. 1) equipped with given reference measures. In particular, our results apply to both categorical and continuous random variables, as the considered manifolds include open subsets of the Euclidean space $\mathbb{R}^{n}$ (equipped with the Lebesgue measure) as well as zero-dimensional discrete spaces $\{1,\dotsc,m\}$ (equipped with the counting measure). For a given pair of smooth manifolds $\mathcal{X}$ and $\mathcal{Y}$, we equip their product $\mathcal{X}\times\mathcal{Y}$ with the product measure and define the set $\mathcal{P}\!\left({\mathcal{X},\mathcal{Y}}\right)$ to consist of all probability measures $P_{XY}$ on $\mathcal{X}\times\mathcal{Y}$ such that the measures $P_{XY}$, $P_{X}(A)=P_{XY}(A\times\mathcal{Y})$ and $P_{Y}(B)=P_{XY}(\mathcal{X}\times B)$ have smooth and positive PDFs (or PMFs) $p_{XY}$, $p_{X}$ and $p_{Y}$ with respect to the reference measures on $\mathcal{X}\times\mathcal{Y}$, $\mathcal{X}$ and $\mathcal{Y}$, respectively. The primary reason for studying $\mathcal{P}\!\left({\mathcal{X},\mathcal{Y}}\right)$ is a convenient formula for pointwise mutual information (Pinsker & Feinstein, 1964, Ch. 2):
Definition 1 (Pointwise Mutual Information).
Let $X$ and $Y$ be random variables valued in $\mathcal{X}$ and $\mathcal{Y}$, respectively, such that $P_{XY}\in\mathcal{P}\!\left({\mathcal{X},\mathcal{Y}}\right)$. The pointwise mutual information (PMI) is defined [^1] We use the natural logarithm, meaning that all quantities are measured in nats. as
$\mathrm{PMI}_{XY}(x,y)=\log\frac{p_{XY}(x,y)}{p_{X}(x)\,p_{Y}(y)}.$
Note that from smoothness and positivity of all the densities it follows that PMI is a smooth, hence measurable, function, s.t. the following distribution exists:
Definition 2 (PMI profile).
The pointwise mutual information profile [^2] Although we are not aware of a prior formal definition and studies of the PMI profile, histograms of approximate PMI between words have been studied before in the computational linguistics community (Allen & Hospedales, 2019). $\mathrm{Prof}_{XY}$ is defined as the distribution of the random variable $T=\mathrm{PMI}_{XY}(X,Y).$
The mutual information is the first moment of the profile, $\mathbf{I}\!\left(X;Y\right)=\mathbb{E}_{T\sim\mathrm{Prof}_{XY}}[T],$ and is known to be invariant under diffeomorphisms (Kraskov et al., 2004, Appendix). More generally, in Appendix A.1 we show that the whole profile is invariant (see Fig. 1):
Theorem 3.
Let $P_{XY}\in\mathcal{P}\!\left({\mathcal{X},\mathcal{Y}}\right)$ and $f\colon\mathcal{X}\to\mathcal{X}$ and $g\colon\mathcal{Y}\to\mathcal{Y}$ be diffeomorphisms. Then for $X^{\prime}=f(X)$ and $Y^{\prime}=g(Y)$ it holds that $P_{X^{\prime}Y^{\prime}}\in\mathcal{P}\!\left({\mathcal{X},\mathcal{Y}}\right)$ and $\mathrm{Prof}_{XY}=\mathrm{Prof}_{X^{\prime}Y^{\prime}}.$
In Appendix A.2 we prove the following three results, characterizing the profiles of simple distributions:
Proposition 4.
Let $X$ and $Y$ be r.v. with joint distribution $P_{XY}\in\mathcal{P}\!\left({\mathcal{X},\mathcal{Y}}\right)$. Then, $\mathbf{I}\!\left(X;Y\right)=0$ if and only if $\mathrm{Prof}_{XY}=\delta_{0}$ is the Dirac measure with a single atom at 00.
Proposition 5.
If $X$ and $Y$ are discrete r.v. with $P_{XY}\in\mathcal{P}\!\left({\mathcal{X},\mathcal{Y}}\right)$, then the PMI profile is discrete:
$\mathrm{Prof}_{XY}=\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}p_{XY}(x,y)\,\delta_{\mathrm{PMI}_{XY}(x,y)}.$
Theorem 6.
Let $X$ and $Y$ be r.v. such that the joint distribution $P_{XY}\in\mathcal{P}\!\left({\mathbb{R}^{m},\mathbb{R}^{n}}\right)$ is multivariate normal. If $k=\min(m,n)$ and $\rho_{1},\rho_{2},\dotsc,\rho_{k}$ are canonical correlations (Jendoubi & Strimmer, 2019) between $X$ and $Y$, then the profile $\mathrm{Prof}_{XY}$ is a generalized $\chi^{2}$ distribution, namely the distribution of the variable
$T=\mathbf{I}\!\left(X;Y\right)+\sum_{i=1}^{k}\frac{\rho_{i}}{2}(Q_{i}-Q^{\prime}_{i}),$
where $Q_{i}$ and $Q^{\prime}_{i}$ are i.i.d. variables sampled according to the $\chi^{2}_{1}$ distribution. In particular, $\mathrm{Prof}_{XY}$ is symmetric around its median, which coincides with the mean $\mathbf{I}\!\left(X;Y\right)=-\frac{1}{2}\sum_{i=1}^{k}\log\left(1-\rho_{i}^{2}\right),$ and all the moments of this distribution exist; its variance is equal to $\rho_{1}^{2}+\cdots+\rho_{k}^{2}$.
However, analytic formulae governing PMI profiles are not known except for the simplest cases above and their reparametrizations by diffeomorphisms due to Theorem 3. Below we show how to construct a family of distributions with diverse PMI profiles, for which the PMI profile can be approximated numerically.
Definition 7.
We call a distribution $P_{XY}\in\mathcal{P}\!\left({\mathcal{X},\mathcal{Y}}\right)$ fine if it is possible to sample $(X,Y)\sim P_{XY}$ and for every point $(x,y)\in\mathcal{X}\times\mathcal{Y}$ we can analytically evaluate the densities $p_{XY}(x,y)$, $p_{X}(x)$ and $p_{Y}(y)$.
For example, discrete distribution as well as multivariate normal and Student distributions are fine. With the advancements in probabilistic computation, we can expect that the list of known fine distributions will grow over time.
The properties are chosen so that we can estimate the PMI profile and mutual information with Monte Carlo approaches. Namely, observe that it is easy to sample $T\sim\mathrm{Prof}_{XY}$ by sampling a data point $(x,y)$ and evaluating $t=\mathrm{PMI}_{XY}(x,y)$. Then, MI can be approximated with a Monte Carlo estimate of the integral $\mathbf{I}\!\left(X;Y\right)=\mathbb{E}[T].$ Assuming $\mathbf{I}\!\left(X;Y\right)<\infty$, the Monte Carlo estimator of the mutual information is guaranteed to be unbiased. For a detailed discussion of Monte Carlo standard error (MCSE) under different regularity conditions see Flegal et al. (2008).
Analogously, to estimate the PMI profile, we can approximate it with a histogram: for bin $B\subset\mathbb{R}$ one can introduce its indicator function $\mathbf{1}_{B}$ and integrate $\mathbb{E}[\mathbf{1}_{B}(T)].$ Its cumulative density function can be approximated with an empirical sample using the expectations $\mathbb{E}[\mathbf{1}_{(-\infty,a_{n}]}(T)]$ for a given sequence $(a_{n})$. As the characteristic functions are bounded between 00 and $1$, the Monte Carlo estimator of both quantities is unbiased and has standard error bounded from above by $1/\sqrt{4n}$ due to the inequality of Popoviciu (1935).
We can enlarge the family of known fine distributions with two constructions, with proofs given in Appendix A.4. First, we can transform known fine distributions with diffeomorphisms with tractable Jacobians, such as normalizing flows (Papamakarios et al., 2021; Kobyzev et al., 2021):
Proposition 8.
If $P_{XY}$ is fine and $f$ and $g$ are diffeomorphisms with tractable Jacobians, then $P_{f(X)g(Y)}$ is fine.
As noted in Theorem 3, the above construction does not change the PMI profile. However, it is possible to obtain more expressive profiles by using finite mixtures of fine distributions (see Fig. 1):
Proposition 9.
Consider a mixture distribution $P_{X^{\prime}Y^{\prime}}$ with weights $w_{1},\dotsc,w_{K}$ and component distributions $P_{X_{1}Y_{1}},\dotsc,P_{X_{K}Y_{K}}$. If all component distributions are fine, then also the mixture distribution is fine.
Since Gaussian mixture models are universal approximators of smooth densities, and they belong to the fine distribution family, in principle one could use these constructions to approximate any distribution. In practice, the evaluation time of PMI grows linearly with the number of components and can become costly for a high number of components (or when large normalizing flow architectures are used). We consider estimating mutual information of mixture distributions an important problem due to their counterintuitive ability to create and destroy information (Haussler & Opper, 1997; Kolchinsky & Tracey, 2017):
Proposition 10.
Consider r.v. $(X_{k},Y_{k})$ such that $\mathbf{I}\!\left(X_{k};Y_{k}\right)<\infty$ for $k=1,\dotsc,K$ and their mixture $(X^{\prime},Y^{\prime})$ with weights $w_{1},\dotsc,w_{K}$. The following inequalities apply:
$0\leq\mathbf{I}\!\left(X^{\prime};Y^{\prime}\right)\leq\log K+\sum_{k=1}^{K}w_{k}\,\mathbf{I}\!\left(X_{k};Y_{k}\right).$
Moreover, these inequalities are tight:
  1. 1.
    There exists a mixture such that $\mathbf{I}\!\left(X^{\prime};Y^{\prime}\right)=\log K$ even though $\mathbf{I}\!\left(X_{k};Y_{k}\right)=0$ for all $k$.
  2. 2.
    There exists a mixture such that $\mathbf{I}\!\left(X^{\prime};Y^{\prime}\right)=0$ even though $\mathbf{I}\!\left(X_{k};Y_{k}\right)>0$ for all $k$.
For completeness, we include a proof in Appendix A.6.
3   Case studies
In this section, we apply fine distributions to three distinct problems. In Sec. 3.1 we demonstrate how they can be used to extend existing benchmarks of mutual information estimators. In Sec. 3.2 we show how fine can be used in experimental sciences to investigate the robustness of mutual information to outliers and inliers. In Sec. 3.3 we apply fine to investigate the biases and sample efficiency of variational estimators employing neural critics.
3.1 Extending the Benchmarks
[…]
In this section, we show how expressive the fine distributions are by visualizing them and demonstrate how they can be applied to investigate the limits of the applicability of mutual information estimators. Czyż et al. (2023) benchmarked mutual information estimators using r.v. $(X,Y)$ distributed according to multivariate normal and Student distributions (for which mutual information is analytically tractable Arellano-Valle et al. (2013)) and their transformations $(f(X),g(Y))$, where $f$ and $g$ are continuous injective functions (which ensures that $\mathbf{I}\!\left(f(X);g(Y)\right)=\mathbf{I}\!\left(X;Y\right)$).
However, transforming only multivariate normal and Student distributions does not ensure that the family of distributions is diverse enough. In particular, if $f$ and $g$ were diffeomorphisms, the diversity of the PMI profiles would be very limited due to Theorem 3. The proposed family of fine distributions can be used as a more expressive alternative to the multivariate normal and Student distributions. Although the analytical formula for ground-truth mutual information of fine distributions is not available, it can be reliably estimated with a Monte Carlo estimator.
To illustrate this application, we implemented four low-dimensional distributions in TensorFlow Probability on JAX (Dillon et al., 2017; Bradbury et al., 2018) assuming that $X$ and $Y$ are valued in Euclidean spaces. Additionally, in Appendix B, we consider cases involving discrete variables. We visualise samples from the distributions considered in Fig. 2 and defer the detailed description to Appendix C.1; the X distribution is a mixture of two bivariate normal distributions. The marginal distributions $P_{X}$ and $P_{Y}$ are normal, although the joint distribution $P_{XY}$ is not. The AI distribution is a mixture of six bivariate normal distributions, illustrating how expressive fine distributions can be. To illustrate that fine distributions are more expressive than mixtures of normal distributions, we consider a two-dimensional $X$ variable and a one-dimensional $Y$ variable. The Waves distribution was created using twelve multivariate normal distributions in which the $X$ variable has been transformed by a diffeomorphism. The Galaxy distribution was created using two multivariate normal distributions in which the $X$ variable was transformed with the spiral diffeomorphism (Czyż et al., 2023).
[…]
Even though the considered problems are low-dimensional and do not encode more information than 1.5 nats, they pose a considerable challenge for the estimators. The performance of the KSG estimator is consistently the best one, with the Waves task not solved by any estimator. The CCA estimator, excelling at distributions that are close to multivariate normal (Czyż et al., 2023), is not able to capture any information at all. This suggests that fine distributions can provide a rich set of distributions that can be used to test mutual information estimators.
3.2 Modeling Outliers
Figure 3: Left: increasing the contamination level $\alpha$ with inlier noise distribution (covariance of $N_{\tilde{Y}}$ is of similar scale as the signal distribution $P_{XY}$). Middle: increasing the contamination level $\alpha$ with outlier noise distribution (noise covariance $N_{\tilde{Y}}$ has $5^{2}$ larger scale than the signal distribution $P_{XY}$). Right: increasing the variance of the noisy normal distribution for constant contamination of $20\%$. Note that outliers (corresponding to very small or large covariance of $n(y)$) have less impact than inliers (corresponding to covariance of $n(y)$ similar to the covariance of the signal distribution $P_{XY}$).
In this section, we use fine distributions to study the effect of inliers and outliers on mutual information estimation. Consider an electric circuit or a biological system modeled as a communication channel $p_{Y\mid X}(y\mid x)$. The researcher controls the input variable $X$, which results in a distribution $P_{X}$, and subsequently measures the outcome variable $Y$. The mutual information $\mathbf{I}\!\left(X;Y\right)$ is then estimated based on the collected experimental samples (Nałȩcz-Jawecki et al., 2023).
However, every experimental system can suffer from occasional failures. We model the output of a failing system with a noise distribution with a PDF $n(y)$. If the probability of system failure does not depend on input value $x$, the communication channel becomes a mixture:
$p_{Y^{\prime}\mid X}(y\mid x)=(1-\alpha)p_{Y\mid X}(y\mid x)+\alpha n(y).$
By multiplying both sides by the density $p_{X}(x)$, the distribution of the channel inputs provided by the scientist, we arrive at the mixture distribution $P_{XY^{\prime}}$ with PDF
$p_{XY^{\prime}}(x,y)=(1-\alpha)p_{XY}(x,y)+\alpha n(y)p_{X}(x).$
If the system failure is unnoticed, one can only measure the r.v. $Y^{\prime}$, rather than $Y$. It is therefore of interest to understand how much $\mathbf{I}\!\left(X;Y\right)$ and $\mathbf{I}\!\left(X;Y^{\prime}\right)$ can differ under realistic assumptions on the noise $n(y)$ and whether the standard mutual information estimation techniques are robust to it. In Appendix A.3 we prove the following upper bound:
Proposition 11.
Consider variables $X$, $Y$ and $Y^{\prime}$, such that
$p_{XY^{\prime}}(x,y)=(1-\alpha)p_{XY}(x,y)+\alpha n(y)p_{X}(x)$
with contamination level $\alpha\in[0,1]$. Then,
$\mathbf{I}\!\left(X;Y^{\prime}\right)\leq(1-\alpha)\,\mathbf{I}\!\left(X;Y\right).$
However, we can use fine distributions to evaluate this quantity exactly. We consider a setting with a two-dimensional input variable $X$ and two-dimensional output variables $Y$ (perfect output) and $Y^{\prime}$ (contaminated output). As the joint density $p_{XY}$ we used a multivariate normal with unit scale and correlations $\mathrm{corr}(X_{1},Y_{1})=\mathrm{corr}(X_{2},Y_{2})=0.8$ and for the noise $n(y)$ we used a multivariate normal distribution with covariance $\sigma^{2}I_{2}$. If $\sigma^{2}\approx 1$ this results in inliers, where the noise distribution is hard to distinguish from the signal. For $\sigma^{2}\ll 1$ the system failures are all close to 00, while outliers are present for $\sigma^{2}\gg 1$.
In Fig. 3 we present the results of three experiments: in the first two, we changed the contamination level $\alpha\in[0,0.5]$ for $\sigma^{2}=1$ (inlier noise) and $\sigma^{2}=5^{2}$ (outlier noise) respectively. In the third experiment, we fixed $\alpha=0.2$ and varied $\sigma^{2}\in\left[2^{-7},2^{8}\right]$. We see that the inlier noise results in a slightly faster decrease of mutual information, while the outlier noise decreases almost linearly following the upper bound $(1-\alpha)\mathbf{I}\!\left(X;Y\right)$. Interestingly, in this low-dimensional setting, the KSG, MINE, and InfoNCE estimators reliably estimate the mutual information $\mathbf{I}\!\left(X;Y^{\prime}\right)$, which can significantly differ from $\mathbf{I}\!\left(X;Y\right)$. Although CCA would be the preferred method to estimate $\mathbf{I}\!\left(X;Y\right)$ without any noise (Czyż et al., 2023), even a small number of outliers ($\alpha=5\%$) can result in unreliable estimates.
3.3 Understanding Neural Critics
Figure 4: Left: PDF of the considered distribution. Middle: neural critic and PMI values. Right: normalized neural critic and PMI profiles.
Figure 5: Estimation of mutual information using a function approximating PMI as a function of sample size for Monte Carlo (MC), InfoNCE, Donsker-Varadhan and NWJ losses. From left to right: true PMI function is used, a constant bias is added, a functional bias is added. The rightmost plot: true PMI function is used in a high-dimensional problem.
Variational estimators of mutual information optimize a critic $f\colon\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$ to obtain an approximate lower bound on mutual information. For example, the Donsker–Varadhan variational approximation (Belghazi et al., 2018)
$I_{\text{DV}}=\sup_{f}\left(\mathbb{E}_{P_{XY}}[f]-\log\mathbb{E}_{P_{X}\otimes P_{Y}}[\exp f]\right)$
is a lower bound and $I_{\text{DV}}=\mathbf{I}\!\left(X;Y\right)$ whenever $f=\mathrm{PMI}_{XY}+c$ for any real number $c$. Analogously, the lower bound of Nguyen et al. (2007),
$I_{\text{NWJ}}=\sup_{f}\left(\mathbb{E}_{P_{XY}}[f]-\mathbb{E}_{P_{X}\otimes P_{Y}}[\exp(f-1)]\right),$
becomes tight for $f=\mathrm{PMI}_{XY}+1$. Oord et al. (2018) propose a variational approximation loss which uses a batch of $(x_{i},y_{i})_{i=1,\dotsc,n}$ samples from $P_{XY}$ to estimate
$I_{\text{NCE}}(f)=\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}\log\frac{\exp f(x_{i},y_{i})}{\frac{1}{n}\sum_{j=1}^{n}\exp f(x_{i},y_{j})}\right]$
Here, if $f(x,y)=\mathrm{PMI}_{XY}(x,y)+c(x),$ where $c$ is any function, then $I_{\text{NCE}}(f)\to\mathbf{I}\!\left(X;Y\right)$ as $n\to\infty$.
In practice, only a finite sample is available, so expectation values are only approximate (and any particular finite-sample estimate does not need to be a lower bound on mutual information) and critic $f$ is modeled via some parametric family, for example, a neural network, which has to be learned from the data available. It is therefore of interest to understand the finite-sample behavior of neural critics and variational approximations.
[…]
How robust are the estimators to a misspecified critic?  In the next experiment (Fig. 5), we evaluated $I_{\text{DV}}$, $I_{\text{NWJ}}$, $I_{\text{InfoNCE}}$, and the simple Monte Carlo estimator (MC) for increasing sample sizes where we provided them with an oracle critic $f$. Using a perfect critic, $f=\mathrm{PMI}_{XY}$, we saw that all estimators had similar performance on this problem. When a constant bias was added, $f=\mathrm{PMI}_{XY}+c$, the Monte Carlo and NWJ estimators became biased. Finally, InfoNCE was the only unbiased estimator when a functional degree of freedom was added, $f(x,y)=\mathrm{PMI}_{XY}(x,y)+\sin(x^{2})$.
In low dimensions, variational approximations do not provide worse mutual information values than the Monte Carlo estimator, when the correct critic function is used, and are more robust to misspecification due to additional degrees of freedom. However, this additional robustness results in a bias in higher dimensions. We simulated a data set using a multivariate normal distribution with 25 strongly interacting components, $\mathrm{corr}(X_{i},Y_{i})=0.8$, which results in $\mathbf{I}\!\left(X;Y\right)\approx 12.8$. Although the Monte Carlo estimator and NWJ estimator (which are not robust to additional degrees of freedom) provided estimates close to the ground-truth (with Monte Carlo slightly outperforming NWJ in both bias and variance), Donsker–Varadhan estimator resulted in a strong positive bias for small batch sizes and InfoNCE had a negative bias with very low variance.
4   Model-based estimation
Figure 6: From left to right: sample from the true data distribution, data distribution sampled using a single MCMC sample, posterior distribution of the mutual information (black line denotes ground-truth value), and posterior distribution of the PMI profiles (black curve denotes the ground-truth value). Top row: well-specified model for the AI distribution. Bottom row: misspecified model for the Galaxy distribution results in biased inferences with miscalibrated posterior. Posterior predictive checking can be used to diagnose unreliable estimates.
As the final application of fine distributions, we consider the problem of model-based estimation of mutual information. Consider a parametric statistical model $\{P_{\theta}\mid\theta\in\Theta\}$ for $(X,Y)$ such that for every value of the parameter vector $\theta$ the mutual information contained in $P_{\theta}$ is known. That is, we define $\mathbf{I}(P_{\theta}):=\mathbf{I}\!\left(X;Y\right)$, where $(X,Y)\sim P_{\theta}$.
If the data are modeled as i.i.d. r.v. $(X_{i},Y_{i})\sim P_{\theta}$, one can estimate $\theta$ and use $\mathbf{I}(P_{\theta})$ as an estimate of the mutual information. For example, Brillinger (2004) interprets the CCA estimator (Kay, 1992) as the estimate $\mathbf{I}(P_{\hat{\theta}})$, where $\hat{\theta}$ is the maximum likelihood estimate in the model in which $P_{\theta}$ is a multivariate normal distribution. Analogously, if $X_{i}$ and $Y_{i}$ are discrete r.v., one can model $\theta$ as a probability matrix and assume $P_{\theta}$ to be the multinomial distribution. Brillinger (2004) proposes to use the maximum likelihood estimate $\hat{\theta}$ also in this case to provide the estimate $\mathbf{I}(P_{\hat{\theta}})$.
More generally, Hutter (2001) observes that if the Dirichlet prior is used for $\theta$, the Bayesian posterior $P(\theta\mid X_{1},Y_{1},\dotsc,X_{N},Y_{N})$ will also be a Dirichlet distribution and provides an approximation via moments to the posterior distribution of mutual information.
We note that one can construct a Bayesian alternative for the CCA estimator by using a prior on the covariance matrix (Lewandowski et al., 2009; Legramanti et al., 2020) and Markov chain Monte Carlo (Gelman et al., 2013, Ch. 11) to provide samples $\theta_{1},\dotsc,\theta_{M}$ from the posterior $P(\theta\mid X_{1},Y_{1},\dotsc,X_{N},Y_{N})$ and construct a sample-based approximation to the posterior on mutual information, $\mathbf{I}(P_{\theta_{1}}),\dotsc,\mathbf{I}(P_{\theta_{M}})$.
More generally, consider a statistical model $\{P_{\theta}\mid\theta\in\Theta\}$ such that all $P_{\theta}$ are fine distributions. One can again find maximum likelihood $\hat{\theta}$ or, if prior $P(\theta)$ is used, apply Bayesian inference algorithms to construct a sample $\theta_{1},\dotsc,\theta_{M}$ from the posterior. Although the exact values for $\mathbf{I}(P_{\hat{\theta}})$ and $\mathbf{I}(P_{\theta_{1}}),\dotsc,\mathbf{I}(P_{\theta_{M}})$ are not available, they can be approximated as in Sec. 2. Since fine distributions include mixture models and normalizing flows, this approach can be used as a general technique for building model-based mutual information estimators, where the generative model can be constructed using domain knowledge. Moreover, this is the first Bayesian estimator of the PMI profile: as all distributions $P_{\theta_{m}}$ are fine, one can construct $M$ histograms (or CDFs) approximating the profile.
To illustrate this approach we implemented a sparse Gaussian mixture model (see Appendix D) in NumPyro (Phan et al., 2019) and used the NUTS sampler (Hoffman & Gelman, 2014) to obtain a Bayesian posterior conditioned on 500 data points from the AI distribution (see Fig. 6). We see that the posterior is concentrated around the ground-truth mutual information value and the ground-truth PMI profile is well-approximated by the posterior samples.
However, this method relies on a model of the true data-generating distribution. As observed in Sec. 3.1, the CCA estimator can yield unreliable estimates when the model is misspecified, that is, the true data-generating process $P_{XY}$ does not belong to the assumed multivariate normal family $P_{\theta}$ (Watson & Holmes, 2016). To illustrate the risks of using misspecified models, we applied the same Gaussian mixture model to 500 points from the Galaxy distribution (Fig. 6). We see that a data sample simulated from the model looks substantially different from the observed data, meaning that the model did not capture the distribution well. In particular, most of the probability mass of the Bayesian posterior is far from the ground-truth mutual information. Similarly, the posterior on the PMI profile is biased. We therefore recommend using model-checking techniques such as posterior predictive checks (Gelman et al., 2013, Ch. 6) and discriminator-based validation (Sankaran & Holmes, 2023, Sec. 4) to understand the deficiencies of the model. To mitigate the risk of overfitting, which can bias the model (see Appendix D), we recommend cross-validation (Piironen & Vehtari, 2017).
5   Conclusion
In this article, we have studied pointwise mutual information profiles, determining them analytically for multivariate normal distributions (Theorem 6), and proposed the family of fine distributions, which include multivariate normal and Student distributions, mixture models and normalizing flows, for which the profile can be approximated using Monte Carlo methods. We showed how fine distributions can be used to provide novel benchmark tasks to test mutual information estimators and calculate mutual information transmitted through a communication channel in the presence of inliers and outliers (which can be used in the experimental design in electrical and biological sciences). Fine distributions allowed us to study how well neural critics approximate pointwise mutual information and we conclude that the error in low-dimensional settings can originate from neural critic misspecification, rather than the formula used to provide a variational approximation. Finally, we showed how Bayesian estimates of mutual information between continuous r.v. can be performed using fine distributions. Additionally, the proposed method provides estimates of the pointwise mutual information profiles, which can be more accurate than the ones obtained using neural critics and variational lower bounds. Although this approach is not universal, we find it suitable for problems with precise domain knowledge available (which can be used to construct the generative distribution $P_{\theta}$ and provide the prior $P(\theta)$) and in which uncertainty quantification is desired.
Limitations and Further Research
In Sec. 3.2, we study a channel in which the probability of failure $\alpha$ is constant and independent of the input variable $X$. However, in both biological and electric systems this assumption may be broken: for large values of $x$ the system may result in an outlier more easily. However, if $\alpha$ depends on $x$, it is not straightforward to use fine distributions to model these situations. Although we propose fine distributions as a basis for model-based Bayesian inference of mutual information, model misspecification (Watson & Holmes, 2016) as well as underfitting and overfitting can bias inference (Piironen & Vehtari, 2017). We therefore recommend following Gelman et al. (2020) to carefully validate the model. Overfitting and model misspecification may be more difficult to detect especially in high-dimensional situations or when more expressive models are constructed, for example using normalizing flows, as Bayesian inference for models involving neural networks is known to be challenging (Izmailov et al., 2021). Finally, fine distributions could be used to study sample size requirements and optimality of different estimation techniques under different distributional assumptions similarly as in Czyż et al. (2023). However, in this work we do not construct a formal benchmark (cf. Sec. 3.1), leaving it for future work. In particular, fine distributions provide a general framework for problems involving both continuous and discrete r.v. and it would be interesting to consider computational examples using other manifolds, for example by using computational geometry frameworks such as geomstats (Miolane et al., 2020).
Acknowledgments
FG was supported by the Norwegian Financial Mechanism GRIEG-1 grant operated by Narodowe Centrum Nauki (National Science Centre, Poland) 2019/34/H/NZ6/00699. PC and AM were supported by a fellowship from the ETH AI Center. We would like to thank Paweł Nałȩcz-Jawecki and Julia Kostin for helpful suggestions on the manuscript.
APPENDIX
Appendix A Technical results
A.1 Proof of the invariance of the pointwise mutual information profile
Recall a well-known result (Kraskov et al., 2004, Appendix):
Lemma 12 (Invariance of PMI).
Let $X^{\prime}=f(X)$ and $Y^{\prime}=g(Y)$, where $f$ and $g$ are diffeomorphisms. Then for every $x^{\prime}$ and $y^{\prime}$ we have
$\mathrm{PMI}_{X^{\prime}Y^{\prime}}(x^{\prime},y^{\prime})=\mathrm{PMI}_{XY}(x,y),$
where $x=f^{-1}(x^{\prime})$ and $y=g^{-1}(y^{\prime})$.
Proof.
From
$p_{X^{\prime}Y^{\prime}}(x^{\prime},y^{\prime})=p_{XY}(x,y)\left|\det D\left(f^{-1}\times g^{-1}\right)(x^{\prime},y^{\prime})\right|$
and analogous quantities we conclude that $p_{X^{\prime}Y^{\prime}}$ as well as $p_{X^{\prime}}$ and $p_{Y^{\prime}}$ are smooth and everywhere positive functions, so that $\mathrm{PMI}_{X^{\prime}Y^{\prime}}$ is well-defined. As $D(f^{-1}\times g^{-1})(x^{\prime},y^{\prime})$ is a block matrix with $Df^{-1}(x^{\prime})$ and $Dg^{-1}(y^{\prime})$ blocks on the diagonal and other blocks zero, we have $\det D(f^{-1}\times g^{-1})(x^{\prime},y^{\prime})=\det Df^{-1}(x^{\prime})\cdot\det Dg^{-1}(y^{\prime})$. ∎
Now we can prove: See 3
Proof.
From the proof of Lemma 12 we conclude that $P_{X^{\prime}Y^{\prime}}\in\mathcal{P}\!\left({\mathcal{X},\mathcal{Y}}\right)$. Then, we note the profile is the pushforward measure
$\mathrm{Prof}_{XY}:=(\mathrm{PMI}_{XY})_{\#}P_{XY}.$
Now let $B\subseteq\mathbb{R}$ be any set in the Borel $\sigma$-algebra and $\mathbf{1}_{B}$ be its characteristic function. Using the change of variables formula for pushforward measure and invariance of PMI:
$\mathrm{Prof}_{X^{\prime}Y^{\prime}}(B)$ $=\int\mathbf{1}_{B}(t)\,\mathrm{d}((\mathrm{PMI}_{X^{\prime}Y^{\prime}})_{\#}P_{X^{\prime}Y^{\prime}})(t)$
$=\int\mathbf{1}_{B}\left(\mathrm{PMI}_{X^{\prime}Y^{\prime}}(x^{\prime},y^{\prime})\right)\,\mathrm{d}P_{X^{\prime}Y^{\prime}}(x^{\prime},y^{\prime})$
$=\int\mathbf{1}_{B}\left(\mathrm{PMI}_{X^{\prime}Y^{\prime}}(f(x),g(y))\right)\,\mathrm{d}P_{XY}(x,y)$
$=\int\mathbf{1}_{B}\left(\mathrm{PMI}_{XY}(x,y)\right)\,\mathrm{d}P_{XY}(x,y)$
$=\mathrm{Prof}_{XY}(B)$
A.2 Pointwise Mutual Information Profiles
The following result shows that the distributions in all $\mathcal{P}\!\left({\mathcal{X},\mathcal{Y}}\right)$ classes with zero mutual information have the same profile:
See 4
Proof.
If $\mathrm{Prof}_{XY}=\delta_{0}$, then the expected value is $\mathbf{I}\!\left(X;Y\right)=0$. To prove the converse, if $\mathbf{I}\!\left(X;Y\right)=0$, then $X$ and $Y$ are independent. Hence, $p_{XY}(x,y)=p_{X}(x)p_{Y}(y)$ at every point $(x,y)$ and $\mathrm{PMI}_{XY}(x,y)=0$ everywhere. ∎
The following result characterizes the PMI profiles for discrete r.v.: See 5
Proof.
The measure $P_{XY}$ is discrete and given by
$P_{XY}=\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}p_{XY}(x,y)\delta_{(x,y)},$
so its pushforward by the $\mathrm{PMI}_{XY}$ function has the form
$\mathrm{Prof}_{XY}=(\mathrm{PMI}_{XY})_{\#}P_{XY}=\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}p_{XY}(x,y)\delta_{\mathrm{PMI}_{XY}(x,y)}.$
See 6
Proof.
Without loss of generality assume that $m\leq n$. As the PMI profile is invariant to diffeomorphisms (Theorem 3), we can also assume that variables $X$ and $Y$ have been whitened by applying canonical correlation analysis (Jendoubi & Strimmer, 2019), that is $\mathbb{E}[X]=0$, $\mathbb{E}[Y]=0$ and the covariance matrix is given by
$\Sigma=\begin{pmatrix}I_{m}&\Sigma_{XY}\\ \Sigma_{XY}^{T}&I_{n}\end{pmatrix}=\begin{pmatrix}I_{m}&R&0\\ R&I_{m}&0\\ 0&0&I_{n-m}\end{pmatrix}$
where
$\Sigma_{XY}=\begin{pmatrix}R&0_{m\times(n-m)}\end{pmatrix}$
is an $m\times n$ matrix with the last $n-m$ columns being zero vectors and $R=\mathrm{diag}(\rho_{1},\dotsc,\rho_{m})$ being the $m\times m$ diagonal matrix representing canonical correlations.
We will write the inverse in the block form
$\Sigma^{-1}=\begin{pmatrix}\Lambda_{X}&\Lambda_{XY}\\ \Lambda_{XY}^{T}&\Lambda_{Y}\end{pmatrix}=\begin{pmatrix}\Lambda_{X}&\tilde{R}&0\\ \tilde{R}&\Lambda_{X}&0\\ 0&0&I_{n-m}\end{pmatrix}$
where the blocks have been calculated using the formula from Petersen & Pedersen (2012, Sec. 9.1):
$\Lambda_{X}$ $=(I_{m}-\Sigma_{XY}\Sigma_{XY}^{T})^{-1}=\mathrm{diag}\left(u_{1},\dotsc,u_{m}\right)$
$\Lambda_{Y}$ $=(I_{n}-\Sigma_{XY}^{T}\Sigma_{XY})=\mathrm{diag}\left(u_{1},\dotsc,u_{m},1,\dotsc,1\right)$
$\Lambda_{XY}$ $=-\Sigma_{XY}\Lambda_{Y}=\begin{pmatrix}\tilde{R}&0_{m\times(n-m)}\end{pmatrix},$
where $\tilde{R}=-\mathrm{diag}\left(u_{1}\rho_{1},\dotsc,u_{m}\rho_{m}\right)$ and $u_{i}=1/\left(1-\rho_{i}^{2}\right)$.
We define a quadratic form
$s(x,y)$ $=x^{T}\Lambda_{X}x+y^{T}\Lambda_{Y}y+2x^{T}\Lambda_{XY}y$
$=\sum_{i=1}^{m}u_{i}\left(x_{i}^{2}+y_{i}^{2}-2\rho_{i}x_{i}y_{i}\right)+\sum_{j=m+1}^{n}y_{j}^{2}$
which can be used to calculate log-PDFs:
$\log p_{XY}(x,y)$ $=-\frac{1}{2}s(x,y)-\frac{1}{2}\log\det\Sigma-\frac{m+n}{2}\log 2\pi,$
$\log p_{X}(x)$ $=-\frac{1}{2}x^{T}x-\frac{m}{2}\log 2\pi,$
$\log p_{Y}(y)$ $=-\frac{1}{2}y^{T}y-\frac{n}{2}\log 2\pi.$
Hence,
$\mathrm{PMI}_{XY}(x,y)=\frac{x^{T}x+y^{T}y-s(x,y)}{2}-\frac{1}{2}\log\det\Sigma.$
We recognize the last summand as
$\mathbf{I}\!\left(X;Y\right)=\frac{1}{2}\log\left(\frac{\det I_{m}\cdot\det I_{n}}{\det\Sigma}\right)=-\frac{1}{2}\log\det\Sigma.$
Define quadratic form
$q(x,y)$ $=2\big{(}\mathrm{PMI}_{XY}(x,y)-\mathbf{I}\!\left(X;Y\right)\big{)}=x^{T}x+y^{T}y-s(x,y)$
$=\sum_{i=1}^{m}\bigg{(}(1-u_{i})\left(x_{i}^{2}+y_{i}^{2}\right)+2\rho_{i}u_{i}x_{i}y_{i}\bigg{)},$
which has a corresponding matrix
$Q=\begin{pmatrix}K&F&0\\ F&K&0\\ 0&0&0\end{pmatrix},$
where
$K=\mathrm{diag}(1-u_{1},\dotsc,1-u_{m})$
and
$F=\mathrm{diag}(\rho_{1}u_{1},\dotsc,\rho_{m}u_{m}).$
We are interested in the distribution of
$q(X,Y)=\begin{pmatrix}X^{T}&Y^{T}\end{pmatrix}Q\begin{pmatrix}X\\ Y\end{pmatrix},$
where $(X,Y)\sim\mathcal{N}(0,\Sigma)$.
Imhof (1961) presents a general approach to evaluating the distributions of such quadratic forms. Consider a r.v.
$Z=\begin{pmatrix}\eta\\ \epsilon\\ \xi\end{pmatrix}\sim\mathcal{N}(0,I_{m+n})$
which is split into blocks of sizes $m$, $m$ and $n-m$. We will construct a linear transformation $A$ such that
$\begin{pmatrix}X\\ Y\end{pmatrix}=A\begin{pmatrix}\eta\\ \epsilon\\ \xi\end{pmatrix}.$
Then, the distribution of $q(X,Y)$ is the distribution of
$Z^{T}A^{T}QAZ,\qquad Z\sim\mathcal{N}(0,I_{m+n}).$
We will construct $A$ as
$A=\begin{pmatrix}P_{-}&P_{+}&0\\ -P_{-}&P_{+}&0\\ 0&0&I_{n-m}\end{pmatrix}$
where
$P_{-}=\mathrm{diag}\left(\sqrt{\frac{1-\rho_{1}}{2}},\cdots,\sqrt{\frac{1-\rho_{m}}{2}}\right),\qquad P_{+}=\mathrm{diag}\left(\sqrt{\frac{1+\rho_{1}}{2}},\cdots,\sqrt{\frac{1+\rho_{m}}{2}}\right).$
We calculate
$AA^{T}=\begin{pmatrix}P_{-}^{2}+P_{+}^{2}&P_{+}^{2}-P_{-}^{2}&0\\ P_{+}^{2}-P_{-}^{2}&P_{-}^{2}+P_{+}^{2}&0\\ 0&0&I_{n-m}\end{pmatrix}=\begin{pmatrix}I_{m}&R&0\\ R&I_{m}&0\\ 0&0&I_{n-m}\end{pmatrix}=\Sigma$
and
$A^{T}QA=\begin{pmatrix}2P_{-}^{2}(K-F)&0&0\\ 0&2P_{+}^{2}(K+F)&0\\ 0&0&0\end{pmatrix},$
where
$2P_{-}^{2}(K-F)=\mathrm{diag}\left(-\rho_{1},\dotsc,-\rho_{m}\right),\qquad 2P_{+}^{2}(K+F)=\mathrm{diag}\left(\rho_{1},\dotsc,\rho_{m}\right).$
Hence, the distribution of $q(X,Y)$ is the same as the distribution of
$\sum_{i=1}^{m}\rho_{i}(-\eta_{i}^{2}+\epsilon_{i}^{2})+\sum_{j=1}^{n-m}0\cdot\xi_{j}^{2},$
where $(\eta_{,}\epsilon,\xi)\sim\mathcal{N}(0,I_{m+n})$. To summarize, let $Q_{1},\dotsc,Q_{m},Q^{\prime}_{1},\dotsc,Q^{\prime}_{m}$ be i.i.d. random variables distributed according to the $\chi^{2}_{1}$ distribution. The quadratic form $q(X,Y)$ has the distribution the same as
$\sum_{i=1}^{m}\rho_{i}(Q_{i}-Q^{\prime}_{i}),$
which can also be written as
$q(X,Y)\sim\sum_{i=1}^{m}\left(\rho_{i}\chi^{2}_{1}-\rho_{i}\chi^{2}_{1}\right).$
Note that this distribution is symmetric around 00. We can now reconstruct the profile from $q(X,Y)$:
$\mathrm{Prof}_{XY}=\mathbf{I}\!\left(X;Y\right)+\sum_{i=1}^{m}\left(\frac{\rho_{i}}{2}\chi^{2}_{1}-\frac{\rho_{i}}{2}\chi^{2}_{1}\right),$
which is symmetric around $\mathbf{I}\!\left(X;Y\right)$ and, in agreement with Proposition 4, degenerates to the atomic distribution $\delta_{0}$ if and only if $\mathbf{I}\!\left(X;Y\right)=0$, which is equivalent to $\rho_{i}=0$ for all $i$.
As a linear combination of independent $\chi^{2}_{1}$ variables, profile has all finite moments. Using the fact that the variance of $\chi^{2}_{1}$ distribution is $2$, and quadratic scaling of variance, each term has variance $22\cdot(\rho_{i}/2)^{2}=\rho_{i}^{2}/$ 2. As variances of independent variables are additive, we can sum up all the $2m$ terms to obtain $\rho_{1}^{2}+\cdots+\rho_{m}^{2}$. ∎
Proposition 13.
With fixed mutual information, the variance of the PMI profile is maximized when $\rho_{1}^{2}=\rho_{2}^{2}=\dotsc=\rho_{m}^{2}$.
Proof.
Let $a_{i}=1-\rho_{i}^{2}$. To maximize the variance we equivalently have to minimize $a_{1}+\dotsc+a_{m}$ preserving given constraint on mutual information and $a_{i}\in(0,1]$.
The constraint on mutual information takes the form
$\mathbf{I}\!\left(X;Y\right)=-\frac{1}{2}\sum_{i=1}^{m}\log\left(1-\rho_{i}^{2}\right)=-\frac{1}{2}\log\left(a_{1}\cdots a_{m}\right).$
Hence, the product $a_{1}\cdots a_{m}$ has to be constant. Denote this constant by $A^{m}$ for $A\in(0,1]$ as well.
Let $a_{1},\dotsc,a_{m}$ be any minimum of $a_{1}+\dotsc+a_{m}$ under the constraints $a_{1}\cdots a_{m}=A^{m}$ and $a_{i}\in(0,1]$. From the inequality between arithmetic and geometric means we note that
$\frac{a_{1}+\dotsc+a_{m}}{m}\geq\sqrt[m]{a_{1}\cdots a_{m}}=A,$
where the equality holds only if $a_{1}=\dotsc=a_{m}=A$. Hence, this is the unique minimum under constraints provided. It follows that $\rho_{1}^{2}=\cdots=\rho_{m}^{2}$. Writing $\rho^{2}$ for the common value, we have
$\rho^{2}=1-\exp\left(-2\mathbf{I}\!\left(X;Y\right)/m\right)$
and
$V=m\rho^{2}=m\left(1-\exp\left(-2\mathbf{I}\!\left(X;Y\right)/m\right)\right).$
The mutual information can also be written as function of variance
$\mathbf{I}\!\left(X;Y\right)=-\frac{1}{2}m\log\left(1-\rho^{2}\right)=-\frac{1}{2}m\log(1-V/m).$
Proposition 14.
With fixed non-zero mutual information the variance of the PMI profile is minimized when $\rho_{i}\neq 0$ for exactly one $i$.
Proof.
Let $a_{i}\in(0,1]$ be any numbers. We have
$(1-a_{1})(1-a_{2})\geq 0$
which is equivalent to
$1+a_{1}a_{2}\geq a_{1}+a_{2}$
where the equality holds if and only if $a_{1}=1$ or $a_{2}=1$.
Using the principle of mathematical induction one can prove a more general inequality:
$a_{1}a_{2}\cdots a_{m}+(m-1)$ $=1+a_{1}(a_{2}\cdots a_{m})+(m-2)$
$\geq a_{1}+\big{(}a_{2}\cdots a_{m}+(m-2)\big{)}$
$\geq a_{1}+a_{2}+\big{(}a_{3}\cdots a_{m}+(m-3)\big{)}$
[…]
$\geq a_{1}+a_{2}+\cdots+a_{m}.$
Let us analyze when the equality can hold. To obtain equality in the first step, we need $a_{1}=1$ or $a_{2}\cdots=a_{m}=1$, which, given the constraints $a_{i}\in(0,1]$ would mean that $a_{2}=\cdots a_{m}=1$. Reasoning inductively, one proves that equality holds only when at least $m-1$ among these numbers are $1$.
Let us apply the above reasoning to the numbers $a_{i}=1-\rho_{i}^{2}$ and note that we are solving a maximization problem $a_{i}\in(0,1]$ under a constraint
$a_{1}\cdots a_{m}=P.$
Using the argument above we note that all the maxima for the above problem are permutations of the sequence $P,1,1,\dotsc,1$. This proves that at most one $\rho_{i}^{2}\neq 0$, what it turns results in at most one $\rho_{i}\neq 0$. ∎
A.3 Proof of the failing channel inequality
In this section we prove:
See 11
Proof.
Let $Z\sim\mathrm{Bernoulli}(1-\alpha)$ be an auxiliary variable. We have $(X,Y^{\prime})\mid Z=1\sim P_{XY}$ and $(X,Y^{\prime})\mid Z=0\sim P_{X}\otimes N_{Y}$.
From the data processing inequality and chain rule we conclude that
$\mathbf{I}\!\left(X;Y^{\prime}\right)\leq\mathbf{I}\!\left(X;Y^{\prime},Z\right)=\mathbf{I}\!\left(X;Z\right)+\mathbf{I}\!\left(X;Y^{\prime}\mid Z\right).$
Now note that $X$ and $Z$ are independent, so $\mathbf{I}\!\left(X;Z\right)=0$.
Hence,
$\mathbf{I}\!\left(X;Y^{\prime}\right)$ $\leq\mathbf{I}\!\left(X;Y^{\prime}\mid Z\right)$
$=\alpha\mathbf{D}_{\mathrm{KL}}\left(P_{XY^{\prime}\mid Z=0}\parallel P_{X\mid Z=0}\otimes P_{Y^{\prime}\mid Z=0}\right)+(1-\alpha)\mathbf{D}_{\mathrm{KL}}\left(P_{XY^{\prime}\mid Z=1}\parallel P_{X\mid Z=1}\otimes P_{Y^{\prime}\mid Z=1}\right)$
$=\alpha\mathbf{D}_{\mathrm{KL}}\left(P_{XY^{\prime}\mid Z=0}\parallel P_{X}\otimes N_{Y}\right)+(1-\alpha)\mathbf{D}_{\mathrm{KL}}\left(P_{XY}\parallel P_{X}\otimes P_{Y}\right)$
$=\alpha\mathbf{D}_{\mathrm{KL}}\left(P_{X}\otimes N_{Y}\parallel P_{X}\otimes N_{Y}\right)+(1-\alpha)\mathbf{I}\!\left(X;Y\right)$
$=(1-\alpha)\mathbf{I}\!\left(X;Y\right).$
A.4 Constructing new fine distributions
In this section we prove Proposition 8 and Proposition 9.
See 8
Proof.
From the proof of Lemma 12 and the assumption of tractability of the Jacobians of $f$ and $g$ we obtain the tractability of the formulae for the densities $p_{f(X)g(Y)}$, $p_{f(X)}$ and $p_{g(Y)}$. Sampling $(f(X),g(Y))$ amounts to sampling $(X,Y)$ and then transforming the sample using $f$ and $g$. ∎
See 9
Proof.
We can evaluate the densities $p_{X^{\prime}Y^{\prime}}(x,y)$, $p_{X^{\prime}}(x)$ and $p_{Y^{\prime}}(y)$ of the mixture distribution using the weighted sums of $p_{X_{k}Y_{k}}(x,y)$, $p_{X_{k}}(x)$ and $p_{Y_{k}}(y)$, respectively. To sample $(X^{\prime},Y^{\prime})\sim P_{X^{\prime}Y^{\prime}}$ from the mixture distribution we can sample an auxiliary variable $Z\sim\mathrm{Categorical}(K;w_{1},\dotsc,w_{K})$. Then, we have $\big{(}(X^{\prime},Y^{\prime})\mid Z=k\big{)}=(X_{k},Y_{k})$, so that we can sample from $P_{X_{k}Y_{k}}$. ∎
A.5 Approximation with smooth densities
In this section, we prove the following result:
Theorem 15.
Let $X$ and $Y$ be random variables with joint distribution $P_{XY}$ and finite mutual information $\mathbf{I}\!\left(X;Y\right)<\infty$. Then, there exists a sequence of random variables $(X_{k},Y_{k})$ for $k=1,2,\dotsc$ such that:
  1. 1.
    Each $P_{X_{k}Y_{k}}\in\mathcal{P}\!\left({m,n}\right)$.
  2. 2.
    Distributions $P_{X_{k}Y_{k}}$ weakly converge to $P_{XY}$ as $k\to\infty$.
  3. 3.
    Mutual information is preserved:
    $\lim_{k\to\infty}\mathbf{I}\!\left(X_{k};Y_{k}\right)=\mathbf{I}\!\left(X;Y\right).$
The proof will automatically follow from two lemmata.
Lemma 16.
Let $X$ be a random variable on $\mathbb{R}^{k}$. Let $N$ be a r.v. independent of $X$ distributed according to the multivariate normal distribution $\mathcal{N}(0,\sigma^{2}I)$ with $\sigma>0$. Then the distribution of $X+N$ has a PDF which is everywhere smooth and positive.
Proof.
Let $P_{X}$ be the distribution of $X$ and $f$ be the PDF of the multivariate normal distribution $\mathcal{N}(0,\sigma^{2}I)$, which is bounded, smooth, and everywhere positive. The convolution of $P_{X}$ and normal density has a PDF
$(f*P_{X})(x)=\int f(x-y)\,\mathrm{d}P_{X}(y)=\mathbb{E}[f(x-X)].$
From the dominated convergence theorem it follows that all the partial derivatives of $\partial_{j}(f*P_{X})$ exist and are given by $(\partial_{j}f)*P_{X}$. From the principle of mathematical induction we conclude that $f*P_{X}$ is smooth.
Positivity follows from the fact that the random variable $f(x-X)$ is strictly positive, so that its expected value is strictly positive as well. ∎
The second lemma is the following:
Lemma 17.
Assume that $I(X;Y)<\infty$. Let $X_{k}=X+N_{k}$ and $Y_{k}=Y+M_{k}$ where $N_{k}$ and $M_{k}$ are independent noise variables distributed according to $\mathcal{N}(0,k^{-2}I)$. Then
$\lim_{k\to\infty}\mathbf{I}\!\left(X_{k};Y_{k}\right)=\mathbf{I}\!\left(X;Y\right).$
Proof.
Observe that for every $k$ the equality $\mathbf{I}\!\left(X_{k};Y_{k}\right)\leq\mathbf{I}\!\left(X;Y\right)$ holds. This is a simple corollary of the data processing inequality applied to the variables $(X_{k},N_{k})$ and $(Y_{k},M_{k})$. Then, note that $P_{X_{k}Y_{k}}$ converges weakly to $P_{XY}$. As mutual information is assumed to be finite, it is given by the Kullback–Leibler divergence, which is lower semicontinuous. This proves the reverse inequality. ∎
A.6 Creating and destroying information with mixtures
Let $A=(0,1)$ and $B=(1,2)$ be two disjoint intervals of unit length. We define two pairs of random variables:
$(X_{1},Y_{1})\sim\mathrm{Uniform}(A\times A),\qquad(X_{2},Y_{2})\sim\mathrm{Uniform}(B\times B).$
Note that
$\mathbf{I}\!\left(X_{1};Y_{1}\right)=\mathbf{I}\!\left(X_{2};Y_{2}\right)=0.$
If $(X,Y)\sim 0.5P_{X_{1}Y_{1}}+0.5P_{X_{2}Y_{2}}$ is distributed according to a mixture, we have
$p_{XY}(x,y)=\frac{1}{2}\mathbf{1}[(x,y)\in A\times A\cup B\times B]$
and
$p_{X}(x)=\frac{1}{2}\mathbf{1}[x\in A\cup B],\qquad p_{Y}(y)=\frac{1}{2}\mathbf{1}[x\in A\cup B].$
Hence,
$\mathrm{PMI}_{XY}(x,y)=\log 2\cdot\mathbf{1}[(x,y)\in A\times A\cup B\times B]$
and
$\mathbf{I}\!\left(X;Y\right)=\frac{1}{2}\log 2+\frac{1}{2}\log 2=\log 2.$
Figure 7: Top row: two distributions with zero MI can be mixed to obtain a distribution with non-zero MI. Bottom row: two distributions with non-zero MI can be mixed to obtain a distribution with zero MI.
For a second example, demonstrating vanishing mutual information, recall the distribution constructed as above:
$p_{XY}(x,y)=\frac{1}{2}\mathbf{1}[(x,y)\in A\times A\cup B\times B]$
and a symmetric one
$p_{UV}(x,y)=\frac{1}{2}\mathbf{1}[(x,y)\in A\times B\cup B\times A].$
We have
$\mathbf{I}\!\left(X;Y\right)=\mathbf{I}\!\left(U;V\right)=\log 2.$
On the other hand, the mixture distribution
$(Z,T)\sim 0.5P_{XY}+0.5P_{UV}$
has zero mutual information, $\mathbf{I}\!\left(Z;T\right)=0$, as
$p_{ZT}(x,y)$ $=\frac{1}{4}\mathbf{1}[(x,y)\in(A\cup B)\times(A\cup B)]$
$=\frac{1}{2}\mathbf{1}[x\in A\cup B]\cdot\frac{1}{2}\mathbf{1}[y\in A\cup B]$
$=p_{Z}(x)\cdot p_{T}(y).$
Note that similar examples can be constructed in the fine distribution family by using mixtures of multivariate normal distributions.
This demonstrates that mixtures can create and destroy information. There is however an upper bound on the amount of information mixtures can create (see Haussler & Opper (1997) and Kolchinsky & Tracey (2017)): See 10
Proof.
The examples provide explicit constructions of distributions with the specified properties. To prove the upper bound, consider a variable $Z\sim\mathrm{Categorical}(K;w_{1},\dotsc,w_{K})$. The random variables corresponding to the mixture distribution, $(X^{\prime},Y^{\prime})$, have conditional distributions
$(X^{\prime},Y^{\prime})\mid Z=k\sim P_{X_{k}Y_{k}}.$
From the data processing inequality and chain rule we have
$\mathbf{I}\!\left(X^{\prime};Y^{\prime}\right)\leq\mathbf{I}\!\left(X^{\prime};Y^{\prime},Z\right)=\mathbf{I}\!\left(X^{\prime};Z\right)+\mathbf{I}\!\left(X^{\prime};Y^{\prime}\mid Z\right).$
As $Z$ is discrete, the first summand, $\mathbf{I}\!\left(X^{\prime};Z\right)$, is bounded from above by the entropy $H(Z)$ (Polyanskiy & Wu, 2022, Th. 3.4(e)), which cannot exceed $\log K$ (Polyanskiy & Wu, 2022, Th. 1.4(b)). The second summand can be written as
$\mathbf{I}\!\left(X^{\prime};Y^{\prime}\mid Z\right)$ $=\sum_{k=1}^{K}P(Z=k)\,\mathbf{D}_{\mathrm{KL}}\left(P_{X^{\prime}Y^{\prime}\mid Z=k}\parallel P_{X^{\prime}\mid Z=k}\otimes P_{Y^{\prime}\mid Z=k}\right)$
$=\sum_{k=1}^{K}w_{k}\,\mathbf{D}_{\mathrm{KL}}\left(P_{X_{k}Y_{k}}\parallel P_{X_{k}}\otimes P_{Y_{k}}\right)=\sum_{k=1}^{K}w_{k}\,\mathbf{I}\!\left(X_{k};Y_{k}\right).$
Appendix B Distributions involving discrete variables
The formalism in Section 2 is applicable to both continuous and discrete random variables, although in Section 3.1 we focus on the distributions in which both $X$ and $Y$ are continuous. If $X$ and $Y$ are discrete, mutual information $\mathbf{I}\!\left(X;Y\right)$ can be calculated analytically from the joint probability matrix and there exist numerous approaches to estimate it from collected samples (Hutter, 2001; Brillinger, 2004).
In this section we consider the mixed case, in which one variable is continuous and the other one is discrete. For example, Carrara & Ernst (2023) describe a particle physics experiment in which $X$ is an 18-dimensional random variable, but $Y$ is binary. Grabowski et al. (2019) consider a cell transmitting information through the MAPK signalling pathway, assuming the input signal $X$ to be discrete and the measured response $Y$ to be continuous.
Yet, there are only a few distributions $P_{XY}$ with known ground-truth mutual information assuming this discrete-continuous case. Gao et al. (2017, Sec. 5) describe a discrete random variable $X$ which is uniformly sampled from the set $\mathcal{X}=\{0,\dotsc,m-1\}$ and the continuous $Y$ variable is sampled as
$(Y\mid X=x)\sim\mathrm{Uniform}(x,x+2),$
which is therefore distributed on $\mathcal{Y}=(0,m+1)$. Gao et al. (2017) prove that mutual information in this case is
$\mathbf{I}\!\left(X;Y\right)=\log m-\frac{m}{m-1}\log 2$
for $m\geq 2$ and $\mathbf{I}\!\left(X;Y\right)=0$ for $m=1$ and consider a multivariate analogue of this distribution, in which pairs of variables $(X_{k},Y_{k})$ for $k=1,\dotsc,K$ are sampled independently using the above procedure. Then, they are concatenated into multivariate vectors $(X_{1},\dotsc,X_{K})$ and $(Y_{1},\dotsc,Y_{K})$ with $K$ times larger mutual information.
Figure 8: Left: samples from the $P_{XY}$ distribution, colored by the value of the binary variable $Y$. Right: the PMI profile of $P_{XY}$ distribution.
We will show how to relate the bivariate example to the framework of fine (multivariate case can be constructed analogously). Note that the joint distribution $P_{XY}$ is not strictly in $\mathcal{P}\!\left({\mathcal{X},\mathcal{Y}}\right)$, as it is supported on the one-dimensional manifold $\mathcal{M}\subset\mathcal{X}\times\mathcal{Y}$ with $m$ connected components (cf. Politis (1991) and Marx et al. (2021)), on which
$p_{XY}(x,y)=\frac{1}{2m}\mathbf{1}[y\in(x,x+2)],$
but it is possible to extend the definitions of Section 2, so that this technical difficulty is resolved. The marginal distributions are tractable and admit PDFs:
$p_{X}(x)$ $=1/m,$
$p_{Y}(y)$ $=\sum_{x=0}^{m-1}p_{XY}(x,y).$
Although $p_{Y}$ is not smooth on $\mathcal{Y}$ (at integer points), this technical difficulty can also be resolved as this set is of measure zero. Hence, although this distribution is not strictly fine, it can still be modelled using the introduced framework.
Next, Gao et al. (2017, Sec. 5) consider the zero-inflated Poissonization of the exponential random variable valued in $\mathcal{X}=\{x\in\mathbb{R}\mid x\geq 0\}$, with: $X\sim\mathrm{Exp}(1)$ and $Y$ being a discrete random variable valued in the set of non-negative integers $\mathcal{Y}=\{0,1,2,\dotsc\}$:
$(Y\mid X=x)\sim p\,\delta_{0}+(1-p)\,\mathrm{Poisson}(x).$
They show that
$\mathbf{I}\!\left(X;Y\right)=(1-p)\left(2\log 2-\gamma-\sum_{k=0}^{\infty}\log k\cdot 2^{-k}\right),$
where $\gamma$ is the Euler–Mascheroni constant.
To use fine distributions in this case, one has to formally extend the definition of $\mathcal{P}\!\left({\mathcal{X},\mathcal{Y}}\right)$ as $\mathcal{X}$ is a manifold with boundary (Lee, 2012, Ch. 1). The joint probability distribution is then given by
$p_{XY}(x,y)=p_{X}(x)\cdot p_{Y\mid X}(y\mid x)=e^{-x}\cdot\left(p\cdot\mathbf{1}[y=0]+(1-p)\frac{e^{-x}x^{y}}{y!}\right)$
and the PMF function of the $Y$ variable is also analytically known:
$p_{Y}(y)=p\cdot\mathbf{1}[y=0]+(1-p)\cdot 2^{-(1+y)}$
Hence, the framework of fine distributions, with minor modifications, can accommodate the above distributions.
However, it also allows one to create more expressive distributions, for which analytical formulae for ground-truth mutual information are not available, but can be approximated with the Monte Carlo methods as explained in Section 2: consider a continuous random variable $X$ and a discrete random variable $Y$. To introduce a dependency between $X$ and $Y$ variable, we can use a mixture of distributions in which the component variables are independent, i.e., $P_{X_{k}Y_{k}}=P_{X_{k}}\otimes P_{Y_{k}}$. Therefore, we consider a graphical model $X\leftarrow Z\rightarrow Y$, in which the distributions $P_{X_{k}}=P_{X\mid Z=k}$ are known and have tractable PDFs. The distributions of $P_{Y_{k}}=P_{Y\mid Z_{k}}$ are given by probability tables. Monte Carlo estimators of Section 2 can then be used to estimate $\mathbf{I}\!\left(X;Y\right)$ with high accuracy. Note also that this general procedure includes the case $X\leftarrow Y$ by setting $Z=Y$.
We illustrate it in a simple example with $\mathcal{X}=\mathbb{R}^{2}$, $\mathcal{Y}=\{0,1\}$ and $K=3$ components.
The first component will model a cluster in the $\mathcal{X}$ space, strongly associated with $Y=1$ value. For $P_{X_{1}}$ we use a bivariate Student distribution centered at $(1,1)$ with isotropic dispersion $\Omega=0.2\cdot I_{2}$ and 8 degrees of freedom. We take $P_{Y_{1}}$ to be the Bernoulli variable with probability $P(Y_{1}=1)=0.95$.
Analogously, we define a second cluster, strongly associated with $Y=0$ value: $P_{X_{2}}$ is a bivariate Student distribution with the same dispersion matrix, but centered at $(-1,-1)$ and with 8 degrees of freedom. Then, $Y_{2}$ is a Bernoulli variable with $P(Y_{2}=1)=0.05$.
We then define a third component using a bivariate normal distribution centered at $(0,0)$ and with covariance matrix
$\Sigma=0.1\begin{pmatrix}1&0.95\\ 0.95&1\end{pmatrix}.$
This component is not informative of $Y$, that is $P(Y_{3}=1)=0.5$.
We used weights $w_{1}=w_{2}=1/4$ and $w_{3}=1/2$, what resulted in the distribution visualised in Fig. 8. We estimated both the profile and the mutual information using $N=10^{6}$ samples and obtained $\mathbf{I}\!\left(X;Y\right)=0.224$ with MCSE of $5.1\cdot 10^{-4}$.
In principle, the above construction can be used to generate realistic high-dimensional data sets (e.g., audio or image) with known ground-truth mutual information, by assuming the generative model $Y\to X$ (i.e., $Z=Y$) and modelling each $P_{X_{k}}$ using a normalizing flow or an autoregressive model (Murphy, 2023, Ch. 22) trained on an auxiliary data set with fixed label $Y_{k}=y_{k}$. Hence, at least in principle, one could obtain highly-expressive generative process $p_{X\mid Y}(x\mid y)$ with tractable probability and sampling. Pairing this with an arbitrary probability vector $p_{Y}(y)$ one can obtain $p_{XY}(x,y)$ and $p_{X}(x)$ even for high-dimensional data sets, so that Monte Carlo estimator can be used to determine the ground-truth mutual information. However, we anticipate possible practical difficulties with scaling up the proposed approach to high-dimensional data and we leave empirical investigation of this topic to future work.
Appendix C Experimental details
C.1 New distributions
We constructed the X distribution as a mixture of bivariate normal distributions with equal weights, zero mean and covariance matrices specified by
$\Sigma_{\pm}=0.3\begin{pmatrix}1&\pm 0.9\\ \pm 0.9&1\end{pmatrix}.$
Note that the marginal distributions of each of component distributions is $\mathcal{N}(0,0.3^{2})$ and subsequently their mixture has exactly the same marginal distributions. This is therefore an interesting example of a distribution in which the joint probability distribution is not multivariate normal, although the marginal distributions of $X$ and $Y$ variables are normal individually.
The AI distribution was constructed as an equally-weighted mixture of six bivariate normal distributions with equal weights and the following parameters:
$\mu_{1}$ $=(1,0)$
$\Sigma_{1}$ $=\mathrm{diag}(0.01,0.2)$
$\mu_{2}$ $=(1,1)$
$\Sigma_{2}$ $=\mathrm{diag}(0.05,0.001)$
$\mu_{3}$ $=(1,-1)$
$\Sigma_{3}$ $=\mathrm{diag}(0.05,0.001)$
$\mu_{4}$ $=(-0.8,-0.2)$
$\Sigma_{4}$ $=\mathrm{diag}(0.03,0.001)$
$\mu_{5}$ $=(-1.2,0)$
$\Sigma_{5}$ $=\begin{pmatrix}0.04&0.085\\ 0.085&0.2\end{pmatrix}$
$\mu_{6}$ $=(-0.4,0)$
$\Sigma_{6}$ $=\begin{pmatrix}0.04&-0.085\\ -0.085&0.2\end{pmatrix}$
The Galaxy distribution was constructed as an equally-weighted mixture of isotropic multivariate normal distributions with $\mu_{\pm}=\pm(1,1,1)$ and unit covariance matrix and the $X$ variable was transformed using the spiral diffeomorphism with $v=0.5$ (cf. Czyż et al. (2023)).
The Waves distribution was created as an equally-weighted mixture of 12 multivariate normal distributions with equal covariance matrices $\Sigma=\mathrm{diag}(0.1,1,0.1)$ and mean vectors
[…]
This construction results in a distribution where different vertical components of the $X$ variable are assigned $Y$ values calculated modulo 4. Then, we transformed the $X$ variable with a continuous injection
$f(x_{1},x_{2})=(x_{1}+5\sin(3x_{2}),x_{2}),$
which does not change the mutual information. Finally, we applied the affine mappings
$a_{1}(x)=0.1x-0.8,\quad a_{2}(y)=0.5y,$
to make the range of the typical values comparable with other distributions.
[…]
C.2 Estimator hyperparameters
Czyż et al. (2023, Appendix E.4) study the effects of hyperparameters on mutual information estimators. We decided to use the histogram-based estimator (Cellucci et al., 2005; Darbellay & Vajda, 1999) with a fixed number of 10 bins per dimension and the popular KSG estimator (Kraskov et al., 2004) with $k=10$ neighbors. Canonical correlation analysis (Kay, 1992; Brillinger, 2004) does not have any hyperparameters. Finally, we employed neural estimators (InfoNCE (Oord et al., 2018) and MINE (Belghazi et al., 2018)) with the neural critic being a ReLU network with 16 and 8 hidden neurons, as it obtained competitive performance in the benchmark of Czyż et al. (2023, Appendix E.4).
As a preprocessing strategy, we followed Czyż et al. (2023, Appendix E.3) and transformed all samples to have zero empirical mean and unit variance along each dimension.
Appendix D Gaussian Mixture Models
Recall from Sec. 4 that for Bayesian estimation of mutual information consists of the following steps:
  1. 1.
    Propose a parametric generative model of the data, $P_{\theta}:=P(X,Y\mid\theta)$, and assume a prior $P(\theta)$ on the parameter space.
  2. 2.
    Use a Markov chain Monte Carlo method to obtain a sample $\theta^{(1)},\dotsc,\theta^{(m)}$ from the posterior $P(\theta\mid X_{1},Y_{1},\dotsc,X_{N},Y_{N})$.
  3. 3.
    Estimate mutual information (and the PMI profile) for each $\theta^{(m)}$ using the Monte Carlo method described in Sec. 2.
  4. 4.
    Validate the findings using e.g., posterior predictive checks and cross-validation.
We consider the following sparse Gaussian mixture model with $K=10$ components:
$\pi$ $\sim\mathrm{Dirichlet}(K;1/K,1/K,\dotsc,1/K),$
$Z_{n}\mid\pi$ $\sim\mathrm{Categorical}(\pi),$ $n=1,\dotsc,N,$
$\mu_{k}$ $\sim\mathcal{N}\left(0,3^{2}I_{D}\right),$ $k=1,\dotsc,K,$
$\Sigma_{k}$ $\sim\mathrm{ScaledLKJ}(1,1),$ $k=1,\dotsc,K,$
$(X_{n},Y_{n})\mid Z_{n},\{\mu_{k},\Sigma_{k}\}$ $\sim\mathcal{N}(\mu_{Z_{n}},\Sigma_{Z_{n}}),$ $n=1,\dotsc,N.$
Sampling a single covariance matrix $\Sigma$ from $\mathrm{ScaledLKJ}(\sigma,\eta)$ distribution corresponds to sampling the correlation matrix $R$ from the Lewandowski-Kurowicka-Joe (LKJ) distribution (Lewandowski et al., 2009):
$p(R)\propto(\det R)^{\eta-1},$
sampling the scale parameters
$\lambda_{1},\lambda_{2},\dotsc,\lambda_{D}\sim\mathrm{HalfCauchy}(\mathrm{scale}{=}\sigma),$
and then constructing the covariance matrix as $\Sigma_{ij}=R_{ij}\lambda_{i}\lambda_{j}$.
The sparse Dirichlet prior is a finite-dimensional alternative to the Dirichlet process, which truncates the number of occupied clusters depending on the data (Frühwirth-Schnatter & Malsiner-Walli, 2019). In particular, the a priori expected number of clusters depends on the number of data points to be observed.
To perform Markov chain Monte Carlo sampling, we implemented the model in NumPyro (Phan et al., 2019), with local latent variables $Z_{n}$ marginalized out and performed sampling using the NUTS sampler (Hoffman & Gelman, 2014). We used 2000 warm-up steps and collected 800 samples.
The $m$th sample is therefore given by
$\theta^{(m)}=\left(\pi^{(m)},\left(\mu_{k}^{(m)},\Sigma_{k}^{(m)}\right)_{k=1,\dotsc,K}\right)$
which is then used to parametrize a Gaussian mixture distribution $P_{\theta^{(m)}}$.
Finally, using Monte Carlo method described in Sec. 2 (with 100,000 samples) we then estimated mutual information $\mathbf{I}(P_{\theta^{(m)}})$ and the profile.
We used the above procedure to perform inference in the model applied to the X, AI, Waves and Galaxy distributions, changing the number of data points $N\in\{125,250,500,1000\}$. We visualise the observed sample, a single posterior predictive sample and posterior on mutual information and the PMI profile in Fig. 9, Fig. 10, Fig. 11 and Fig. 12.
Figure 9: Gaussian mixture model fitted to the X distribution with 125, 250, 500 and 1000 samples.
Figure 10: Gaussian mixture model fitted to the AI distribution with 125, 250, 500 and 1000 samples.
Figure 11: Gaussian mixture model fitted to the Waves distribution with 125, 250, 500 and 1000 samples.
Figure 12: Gaussian mixture model fitted to the Galaxy distribution with 125, 250, 500 and 1000 samples.